OpenStudy (anonymous):

what is the density of uranium hexaflouride at 127C and 763mmHg?

OpenStudy (jfraser):

use and rearrange the ideal gas law to solve for the \(density\) of a gas. Since:\[P*V = n*R*T\]and\[n = \frac{mass}{MM}\]then\[P*V = \frac{mass*R*T}{MM}\]so \[density = \frac{mass}{V} = \frac{MM*P}{R*T}\]

OpenStudy (jfraser):

where \(MM\) is the molar mass